# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4 Realisations of the cyclotomic Gaudin model
In this section, we study two different realisations of the cyclotomic Gaudin model – the periodic Toda chain and the discrete self-trapping (DST) model – with the objective of describing their corresponding hierarchies variationally. We will then go on to show how our framework allows for a straightforward coupling of integrable hierarchies, by using the approach devised in [CSV, Section 7] to couple together hierarchies of integrable field theories.
In what follows, we will work with the cyclotomic Gaudin model associated with the Lie algebra $\mathfrak{g}\coloneqq\mathfrak{gl}_{T}(\mathbb{C})$ and the automorphism $\sigma\in\text{Aut}\,\mathfrak{g}$ defined by $\sigma(E_{ij})=\omega^{j-i}E_{ij}$, for every $i,j=1,\ldots,T$. Here $\omega$ is a primitive $T$th root of unity, and by $E_{ij}$, $i,j=1,\ldots,T$, we denote the standard basis of $\mathfrak{gl}_{T}(\mathbb{C})$ taking the indices $i$ and $j$ modulo $T$ by convention. The eigenspaces of $\sigma$ defined by (3.3) are then given by $\mathfrak{g}^{(n)}=\text{span}\{E_{i,i+n}\}_{i=1}^{T}$.
Let us also note the following useful identity that we shall frequently make use of:
$\frac{z^{T-1-[l]}w^{[l]}}{z^{T}-w^{T}}=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\omega^{-kl}}{z-\omega^{k}w}$ (4.1)
for any $z,w\in\mathbb{C}$ and $l\in\mathbb{Z}$, where $[l]\coloneqq l$ mod $T$.
@@ 4.1 Periodic Toda chain
The periodic Toda chain [T] describes a system of particles connected by “exponential springs” together with a periodic boundary condition, and has been extensively studied in the Hamiltonian formalism. See, for instance, [F] for the widely used Flaschka change of coordinates, and [AM] for a proof of its integrability.
Lax matrix:
We will work with the Lax matrix
$L_{\text{Toda}}(\lambda)=\left(\begin{matrix}\leavevmode\nobreak\ p_{1}\lambda^{-1}\leavevmode\nobreak\ &1&0&&\dots&&a_{T}\lambda^{-2}\\ a_{1}\lambda^{-2}&p_{2}\lambda^{-1}&1&&\dots&&0&\\[4.30554pt] \vdots&&\ddots&&&&\vdots\\ 0&&a_{i-1}\lambda^{-2}&p_{i}\lambda^{-1}&1&&0\\ \vdots&&&&\ddots&&\vdots\\ 0&&\dots&&a_{T-2}\lambda^{-2}&p_{T-1}\lambda^{-1}&1\\ 1&&\dots&&0&a_{T-1}\lambda^{-2}&\leavevmode\nobreak\ p_{T}\lambda^{-1}\end{matrix}\!\!\!\!\right)$ (4.2)
where $a_{i}=e^{q_{i}-q_{i+1}}$, and $q_{i},p_{i}$ are the canonical coordinates satisfying the canonical Poisson bracket relations $\{p_{i},q_{j}\}=\delta_{ij}$, for $i,j=1,\ldots,T$. We also have the periodic boundary conditions $(p_{0},q_{0})=(p_{T},q_{T})$ and $(p_{T+1},q_{T+1})=(p_{1},q_{1})$.
The standard Lax matrix for the periodic Toda chain (see [BBT, Chapter 6], for instance)
$\widetilde{L}_{\text{Toda}}(\lambda)=\left(\;\begin{matrix}\leavevmode\nobreak\ p_{1}\leavevmode\nobreak\ &a_{1}^{1/2}&0&&\dots&&a_{T}^{1/2}\lambda^{-1}\\ a_{1}^{1/2}&p_{2}&a_{2}^{1/2}&&\dots&&0&\\[4.30554pt] \vdots&&\ddots&&&&\vdots\\ 0&&a_{i-1}^{1/2}&p_{i}&a_{i}^{1/2}&&0\\ \vdots&&&&\ddots&&\vdots\\ 0&&\dots&&a_{T-2}^{1/2}&p_{T-1}&a_{T-1}^{1/2}\\ a_{T}^{1/2}\lambda&&\dots&&0&a_{T-1}^{1/2}&\leavevmode\nobreak\ p_{T}\leavevmode\nobreak\ \\ \end{matrix}\!\!\!\!\right)$ (4.3)
is related to $L_{\text{Toda}}(\lambda)$ by conjugation by the diagonal matrix $\mathcal{Q}=\text{diag}(e^{-q_{1}/2}\lambda^{-1},\ldots,e^{-q_{T}/2}\lambda^{-T})$ and multiplication by an overall factor of $\lambda^{-1}$, together with a change of $\lambda$-dependence, as follows
$L_{\text{Toda}}(\lambda)=\lambda^{-1}\mathcal{Q}\widetilde{L}_{\text{Toda}}(\lambda^{T})\mathcal{Q}^{-1}\,.$ (4.4)
@@
$\big{\{}\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda),\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\big{\}}=\big{[}\widetilde{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda)+\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\big{]}$ (4.5)
@@
$\widetilde{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{2}\frac{\mu+\lambda}{\mu-\lambda}\sum_{i=1}^{T}E_{ii}\otimes E_{ii}+\frac{1}{\mu-\lambda}\bigg{(}\mu\sum_{i<j}+\lambda\sum_{i>j}\bigg{)}E_{ij}\otimes E_{ji}\,.$ (4.6)
@@
$\{L_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda),L_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\}=[r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),L_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda)]-[r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),L_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)]$ (4.7)
@@
$r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\omega^{k(j-i)}}{\mu-\omega^{-k}\lambda}E_{ij}\otimes E_{ji}\,.$ (4.8)
that we have been working with. This explains our choice of Lax matrix (4.2) as opposed to the more traditional one. As just proved, it satisfies the the Poisson algebra of the cyclotomic $\mathfrak{gl}_{T}(\mathbb{C})$-Gaudin model and therefore allows us to obtain the periodic Toda chain as a certain realisation of that model. Specifically, the Lax matrix $L_{\text{Toda}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix with double poles at the origin and at infinity, that is,
$L_{\text{Toda}}(\lambda)=\frac{J_{0}^{(0)}}{\lambda}+\frac{J_{0}^{(1)}}{\lambda^{2}}+J_{\infty}$ (4.9)
@@
$J_{0}^{(0)}=\sum_{i=1}^{T}p_{i}E_{ii}\,,\quad J_{0}^{(1)}=\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}E_{i+1,i}\,,\quad J_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\,.$ (4.10)
Orbit realisation:
Set $D=\{0,\infty\}$, and choose the non-dynamical element
$\Lambda_{\text{Toda}}(\lambda)=\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\Lambda_{\infty}\in\Omega_{D}^{\Gamma}$ (4.11)
@@
$\Lambda_{0}^{(1)}=\sum_{i=1}^{T}E_{i+1,i}\in\mathfrak{g}^{(-1)}\,,\quad\Lambda_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\in\mathfrak{g}^{(1)}\,.$ (4.12)
@@
$J_{0}^{(0)}=[\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},J_{0}^{(1)}]\,,\quad J_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1}\,,\quad J_{\infty}=\Lambda_{\infty}\,.$ (4.13)
@@
$J_{0}^{(0)}=\sum_{i=1}^{T}\left(\frac{v_{i}}{u_{i}}-\frac{v_{i-1}}{u_{i-1}}\right)E_{ii}\,,\quad J_{0}^{(1)}=\sum_{i=1}^{T}\dfrac{u_{i+1}}{u_{i}}E_{i+1,i}\,,\quad J_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\,.$ (4.15)
@@
$p_{i}=\frac{v_{i}}{u_{i}}-\frac{v_{i-1}}{u_{i-1}}\quad\text{and}\quad q_{i}=-\ln{u_{i}}\,,\quad\text{for}\,\,i=1,\ldots,T\,,$ (4.16)
we now have the desired realisation of the coefficients of the Lax matrix $L_{\text{Toda}}(\lambda)$. The coadjoint orbit ${\cal O}_{\Lambda}^{\text{Toda}}$ where $L_{\text{Toda}}(\lambda)$ lives is parameterised by $a_{i},p_{i}$, $i=1,\ldots,T$, satisfying $\prod_{i}^{T}a_{i}=1$ and $\sum_{i}^{T}p_{i}=0$.
Lax equations:
Let us now look at the Lax equations associated with the invariant functions on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ defined in (3.70) for the general case. The only functions we need to consider are
$H_{p,0}=\dfrac{1}{p+1}\operatorname{Res}_{\lambda=0}(\lambda^{p}\operatorname{Tr}(\iota_{\lambda_{0}}L_{\text{Toda}}^{p+1}))d\lambda\,.$ (4.17)
@@
$\partial_{t_{p}^{0}}\iota_{\lambda}L_{\text{Toda}}=[R_{\pm}\nabla H_{p,0}(\iota_{\lambda}L_{\text{Toda}}),\iota_{\lambda}L_{\text{Toda}}]\,.$ (4.18)
@@
$\partial_{t_{1}^{0}}L_{\text{Toda}}=[M_{1,0},L_{\text{Toda}}]\quad\text{with}\quad M_{1,0}=-\dfrac{J_{0}^{(1)}}{\lambda}\,.$ (4.19)
Taking residues on both sides in (4.19) gives the equations of motion for $p_{i},q_{i}$, for $i=1,\ldots,T$. To get the equations of motion for $q_{i}$ it is most convenient to “undo” the dressing and write the corresponding equation as
$[\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)},J_{0}^{(1)}]=0\,.$ (4.20)
This tells us that the diagonal matrix $\phi_{0}^{(0)\,-1}\left(\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)}\right)\phi_{0}^{(0)}$ must commute with $\Lambda_{0}^{(1)}$, and is therefore equal to $\alpha\bm{1}$. Using the freedom $\phi_{0}^{(0)}\to\phi_{0}^{(0)}g$, where $g$ is diagonal, to set $\det\phi_{0}^{(0)}=1$, we see that $\alpha=0$ by taking the trace of $\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)}=\alpha\bm{1}$. Thus, we have the desired Toda equations
$\begin{split}&\partial_{t_{1}^{0}}q_{i}=-p_{i}\,,\\ &\partial_{t_{1}^{0}}p_{i}=e^{q_{i}-q_{i+1}}-e^{q_{i-1}-q_{i}}\,.\end{split}$ (4.21)
Lagrangian description:
Using Theorem 3.6, we can now write a Lagrangian multiform for the periodic Toda hierarchy as
$\mathscr{L}_{\text{Toda}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\mathrm{d}t_{p}^{0}$ (4.22)
@@
$\mathscr{L}_{p,0}=-\sum_{i=1}^{T}p_{i}\partial_{t_{p}^{0}}q_{i}-\dfrac{1}{p+1}\operatorname{Res}_{\lambda=0}(\lambda^{p}\operatorname{Tr}(L_{\text{Toda}}^{p+1}))d\lambda\,.$ (4.23)
@@
$\mathscr{L}_{1,0}=-\sum_{i=1}^{T}p_{i}\partial_{t_{1}^{0}}q_{i}-\frac{1}{2}\sum_{i=1}^{T}p_{i}^{2}-\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}\,,$ (4.24)
This is the expected phase space Lagrangian that our method produces and which gives Hamilton’s equations (4.21) for periodic Toda. It corresponds to the (tangent bundle) Lagrangian $\mathscr{L}_{1,0}=\frac{1}{2}\sum_{i=1}^{T}(\partial_{t_{1}^{0}}q_{i})^{2}-\sum_{i=1}^{T}e^{q_{i}-q_{i+1}}$.

### v2

@@ 4 Realisations of the cyclotomic Gaudin model
In this section, we study two different realisations of the cyclotomic Gaudin model – the periodic Toda chain and the discrete self-trapping (DST) model – with the objective of describing their corresponding hierarchies variationally. We will then go on to show how our framework allows for a straightforward coupling of integrable hierarchies, by using the approach devised in [15, Section 7] to couple together hierarchies of integrable field theories.
In what follows, we will work with the cyclotomic Gaudin model associated with the Lie algebra $\mathfrak{g}\coloneqq\mathfrak{gl}_{T}(\mathbb{C})$ and the automorphism $\sigma\in\text{Aut}\,\mathfrak{g}$ defined by $\sigma(E_{ij})=\omega^{j-i}E_{ij}$, for every $i,j=1,\dots,T$. Here $\omega$ is a primitive $T$-th root of unity, and by $E_{ij}$, $i,j=1,\dots,T$, we denote the standard basis of $\mathfrak{gl}_{T}(\mathbb{C})$ taking the indices $i$ and $j$ modulo $T$ by convention. The eigenspaces of $\sigma$ defined by (3.1) are then given by $\mathfrak{g}^{(n)}=\text{span}\{E_{i,i+n}\}_{i=1}^{T}$.
Let us also note the following useful identity that we shall frequently make use of
$\frac{z_{1}^{T-1-[l]}z_{2}^{[l]}}{z_{1}^{T}-z_{2}^{T}}=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\omega^{-kl}}{z_{1}-\omega^{k}z_{2}}$ (4.1)
for any $z_{1},z_{2}\in\mathbb{C}$ and $l\in\mathbb{Z}$, where $[l]\in\{0,\dots,T-1\}$ is such that $l=[l]\mod T$.
@@ 4.1 Periodic Toda chain
The periodic Toda chain [39] describes a system of particles connected by “exponential springs” together with a periodic boundary condition, and has been extensively studied in the Hamiltonian formalism. See, for instance, [18] for the widely used Flaschka change of coordinates, and [3] for a proof of its integrability.
Lax matrix. We will work with the Lax matrix
$L_{\text{Toda}}(\lambda)=\begin{pmatrix}p_{1}\lambda^{-1}&1&0&&\dots&&a_{T}\lambda^{-2}\\ a_{1}\lambda^{-2}&p_{2}\lambda^{-1}&1&&\dots&&0&\\[4.30554pt] \vdots&&\ddots&&&&\vdots\\ 0&&a_{i-1}\lambda^{-2}&p_{i}\lambda^{-1}&1&&0\\ \vdots&&&&\ddots&&\vdots\\ 0&&\dots&&a_{T-2}\lambda^{-2}&p_{T-1}\lambda^{-1}&1\\ 1&&\dots&&0&a_{T-1}\lambda^{-2}&p_{T}\lambda^{-1}\end{pmatrix},$ (4.2)
where $a_{i}={\rm e}^{q_{i}-q_{i+1}}$, and $q_{i}$, $p_{i}$ are the canonical coordinates satisfying the canonical Poisson bracket relations $\{p_{i},q_{j}\}=\delta_{ij}$ for $i,j=1,\dots,T$. We also have the periodic boundary conditions $(p_{0},q_{0})=(p_{T},q_{T})$ and $(p_{T+1},q_{T+1})=(p_{1},q_{1})$.
The standard Lax matrix for the periodic Toda chain (see [7, Section 6], for instance)
$\widetilde{L}_{\text{Toda}}(\lambda)=\begin{pmatrix}p_{1}&a_{1}^{1/2}&0&&\dots&&a_{T}^{1/2}\lambda^{-1}\\ a_{1}^{1/2}&p_{2}&a_{2}^{1/2}&&\dots&&0&\\[4.30554pt] \vdots&&\ddots&&&&\vdots\\ 0&&a_{i-1}^{1/2}&p_{i}&a_{i}^{1/2}&&0\\ \vdots&&&&\ddots&&\vdots\\ 0&&\dots&&a_{T-2}^{1/2}&p_{T-1}&a_{T-1}^{1/2}\\ a_{T}^{1/2}\lambda&&\dots&&0&a_{T-1}^{1/2}&p_{T}\\ \end{pmatrix}$
is related to $L_{\text{Toda}}(\lambda)$ by conjugation by the diagonal matrix $\mathcal{Q}\!=\!{\rm diag}\bigl{(}{\rm e}^{-q_{1}/2}\lambda^{-1},\dots,{\rm e}^{-q_{T}/2}\lambda^{-T}\bigr{)}$ and multiplication by an overall factor of $\lambda^{-1}$, together with a change of $\lambda$-dependence, as follows:
$L_{\text{Toda}}(\lambda)=\lambda^{-1}\mathcal{Q}\widetilde{L}_{\text{Toda}}\bigl{(}\lambda^{T}\bigr{)}\mathcal{Q}^{-1}.$ (4.3)
@@
$\bigl{\{}\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda),\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{\}}=\bigl{[}\widetilde{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda)+\widetilde{L}_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{]},$
@@
$\widetilde{r}_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{2}\frac{\mu+\lambda}{\mu-\lambda}\sum_{i=1}^{T}E_{ii}\otimes E_{ii}+\frac{1}{\mu-\lambda}\biggl{(}\mu\sum_{i<j}+\lambda\sum_{i>j}\biggr{)}E_{ij}\otimes E_{ji}.$
@@
$\bigl{\{}L_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda),L_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{\}}=\bigl{[}r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu),L_{\text{Toda}\,_{\mathbf{\underline{1}}}}(\lambda)\bigr{]}-\bigl{[}r_{{}_{\mathbf{\underline{21}}}}(\mu,\lambda),L_{\text{Toda}\,_{\mathbf{\underline{2}}}}(\mu)\bigr{]},$
@@
$r_{{}_{\mathbf{\underline{12}}}}(\lambda,\mu)=\frac{1}{T}\sum_{k=0}^{T-1}\frac{\omega^{k(j-i)}}{\mu-\omega^{-k}\lambda}E_{ij}\otimes E_{ji}\,$ (4.4)
that we have been working with. This explains our choice of Lax matrix (4.2) as opposed to the more traditional one. As just proved, it satisfies the Poisson algebra of the cyclotomic $\mathfrak{gl}_{T}(\mathbb{C})$-Gaudin model and therefore allows us to obtain the periodic Toda chain as a certain realisation of that model. Specifically, the Lax matrix $L_{\text{Toda}}(\lambda)$ can be seen as a realisation of the cyclotomic Gaudin Lax matrix with double poles at the origin and at infinity, that is,
$L_{\text{Toda}}(\lambda)=\frac{J_{0}^{(0)}}{\lambda}+\frac{J_{0}^{(1)}}{\lambda^{2}}+J_{\infty},$ (4.5)
@@
$J_{0}^{(0)}=\sum_{i=1}^{T}p_{i}E_{ii},\qquad J_{0}^{(1)}=\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}E_{i+1,i},\qquad J_{\infty}=\sum_{i=1}^{T}E_{i,i+1}.$
Orbit realisation. Set $D=\{0,\infty\}$, and choose the non-dynamical element
$\Lambda_{\text{Toda}}(\lambda)=\frac{\Lambda_{0}^{(1)}}{\lambda^{2}}+\Lambda_{\infty}\in\Omega_{D^{\prime}}^{\Gamma},$
@@
$\Lambda_{0}^{(1)}=\sum_{i=1}^{T}E_{i+1,i}\in\mathfrak{g}^{(-1)},\qquad\Lambda_{\infty}=\sum_{i=1}^{T}E_{i,i+1}\in\mathfrak{g}^{(1)}.$
@@
$J_{0}^{(0)}=\bigl{[}\phi_{0}^{(1)}\phi_{0}^{(0)\,-1},J_{0}^{(1)}\bigr{]},\qquad J_{0}^{(1)}=\phi_{0}^{(0)}\Lambda_{0}^{(1)}\phi_{0}^{(0)\,-1},\qquad J_{\infty}=\Lambda_{\infty}.$ (4.6)
@@
$J_{0}^{(0)}=\sum_{i=1}^{T}\biggl{(}\frac{v_{i}}{u_{i}}-\frac{v_{i-1}}{u_{i-1}}\biggr{)}E_{ii},\qquad J_{0}^{(1)}=\sum_{i=1}^{T}\dfrac{u_{i+1}}{u_{i}}E_{i+1,i},\qquad J_{\infty}=\sum_{i=1}^{T}E_{i,i+1}.$
@@
$p_{i}=\frac{v_{i}}{u_{i}}-\frac{v_{i-1}}{u_{i-1}}\qquad\text{and}\qquad q_{i}=-\ln{u_{i}}\qquad\text{for}\quad i=1,\dots,T,$
we now have the desired realisation of the coefficients of the Lax matrix $L_{\text{Toda}}(\lambda)$. The coadjoint orbit ${\cal O}_{\Lambda}^{\text{Toda}}$ where $L_{\text{Toda}}(\lambda)$ lives is parameterised by $a_{i}$, $p_{i}$, $i=1,\dots,T$, satisfying $\prod_{i}^{T}a_{i}=1$ and $\sum_{i}^{T}p_{i}=0$.
Lax equations. Let us now look at the Lax equations associated with the invariant functions on $\widetilde{\mathfrak{g}}_{D}^{(1)}$ defined in (3.36) for the general case. The only functions we need to consider are
$H_{p,0}=\dfrac{1}{p+1}\operatorname{Res}_{\lambda=0}\bigl{(}\lambda^{p}\operatorname{Tr}\bigl{(}\iota_{\lambda_{0}}L_{\text{Toda}}^{p+1}\bigr{)}\bigr{)}{\rm d}\lambda.$
@@
$\partial_{t_{p}^{0}}\iota_{\lambda}L_{\text{Toda}}=[R_{\pm}\nabla H_{p,0}(\iota_{\lambda}L_{\text{Toda}}),\iota_{\lambda}L_{\text{Toda}}].$
@@
$\partial_{t_{1}^{0}}L_{\text{Toda}}=[M_{1,0},L_{\text{Toda}}]\qquad\text{with}\quad M_{1,0}=-\dfrac{J_{0}^{(1)}}{\lambda}.$ (4.7)
Taking residues on both sides in (4.7) gives the equations of motion for $p_{i}$, $q_{i}$ for $i=1,\dots,T$. To get the equations of motion for $q_{i}$ it is most convenient to “undo” the dressing and write the corresponding equation as
$\bigl{[}\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)},J_{0}^{(1)}\bigr{]}=0.$
This tells us that the diagonal matrix
$\phi_{0}^{(0)\,-1}\bigl{(}\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)}\bigr{)}\phi_{0}^{(0)}$
must commute with $\Lambda_{0}^{(1)}$, and is therefore equal to $\alpha\bm{1}$. Using the freedom $\phi_{0}^{(0)}\to\phi_{0}^{(0)}g$, where $g$ is diagonal, to set $\det\phi_{0}^{(0)}=1$, we see that $\alpha=0$ by taking the trace of
$\partial_{t_{1}^{0}}\phi_{0}^{(0)}\phi_{0}^{(0)\,-1}-J_{0}^{(0)}=\alpha\bm{1}.$
Thus, we have the desired Toda equations
$\partial_{t_{1}^{0}}q_{i}=-p_{i},$
$\partial_{t_{1}^{0}}p_{i}={\rm e}^{q_{i}-q_{i+1}}-{\rm e}^{q_{i-1}-q_{i}}.$ (4.8)
Lagrangian description. Using Theorem 3.6, we can now write a Lagrangian multiform for the periodic Toda hierarchy as
$\mathscr{L}_{\text{Toda}}=\sum_{p=1}^{M}\mathscr{L}_{p,0}\,\mathrm{d}t_{p}^{0}$
@@
$\mathscr{L}_{p,0}=-\sum_{i=1}^{T}p_{i}\partial_{t_{p}^{0}}q_{i}-\dfrac{1}{p+1}\operatorname{Res}_{\lambda=0}\bigl{(}\lambda^{p}\operatorname{Tr}\bigl{(}L_{\text{Toda}}^{p+1}\bigr{)}\bigr{)}{\rm d}\lambda.$
@@
$\mathscr{L}_{1,0}=-\sum_{i=1}^{T}p_{i}\partial_{t_{1}^{0}}q_{i}-\frac{1}{2}\sum_{i=1}^{T}p_{i}^{2}-\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}.$
This is the expected phase space Lagrangian that our method produces and which gives Hamilton’s equations (4.8) for periodic Toda. It corresponds to the (tangent bundle) Lagrangian
$\mathscr{L}_{1,0}=\frac{1}{2}\sum_{i=1}^{T}\bigl{(}\partial_{t_{1}^{0}}q_{i}\bigr{)}^{2}-\sum_{i=1}^{T}{\rm e}^{q_{i}-q_{i+1}}.$